THE OPTICAL STARK SPECTRUM OF the [17.8]0^{+}−X^{1}Σ^{+} BAND OF GOLD MONOFLUORIDE,AuF

RUOHAN ZHANG, CHENGBING QIN, AND \underline{TIMOTHY C. STEIMLE}, Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287; THOMAS VARBERG, Department of Chemistry,Macalaster College,1600 Grand Avenue, St. Paul,MN 55105.

The bonding and electrostatic prperties of gold containig molecules are highly influenced by realtivistic effects^{a}. Recently Varberg’s group reported^{b} on the field-free detection and analysis of the [17.8]0^{+}−X^{1}Σ^{+} band system of AuF at Doppler limited resolution. Here we report on the analysis of the same band recorded field free and in the presence of a static electric field for a cold molecular beam sample. The Stark induced shifts of the R(0)(17756.156 cm^{−1}), R(1)(17756.640 cm^{−1}) and P(1)(17755.123 cm^{−1})branch features were analyzed to determine the permanent electric dipole moment, μ_{el}, of 2.025 D and 4.218 D for the [17.8]0^{+} and X^{1}Σ^{+} states, respectively. The agreement with the various DFT ^{c} and wave-function^{d} based predictions will be discussed.

aP.Pyykko; Angew Chem. Int 43 4412, 2004.

bE.K. Butler, B. J. Knurr, K. J. Manke, T. R. Vervoot, and T. D. Varberg; J. Phys. Chem. A 114 4831, 2010.

cE. Goll, and H. Stoll; Phys. Rev. A 76 032507, 2007.

dF. Holka, A. Avramopoulos, O. Loboda, V. Kello, and M.G. Papadopoulos; Chem.Phys. Lett 472 185, 2009.